find each measure of the circle

A Find each measure. 26. MLAPD 28. mABD / 1600B 27. mAC 29. mZCPA L
Answer #1

a solution :- Im LAPD = 180-90-60 - (30° 27 mac = 90+30 = 120° [28 mABD 60+ 180 +90 : c[330] 39 mLCPA = 90+30 = 120